# Sea ice floe segmentation in close-range optical imagery using active contour and foundation models

### v4

@@ 5.2.2 Diffuse sea ice boundaries
Of the various methods, SAM-P provides the best overall balance of accuracy and computational efficiency across both the distinct edges and diffuse boundaries images. SAM-P has the fewest missed floes and maintains stable, though slightly higher, rates of misinterpretations and hallucinations compared to SAM-GVF. While SAM-GVF performs marginally better in terms of misinterpretation and hallucination metrics, these improvements come at a substantial cost in computational time ($\approx$15 times slower per floe). SAM-A and GVF show significant deterioration, particularly with a substantial increase in missed floes. Additionally, computational time per floe increases for GVF (and consequently for SAM-GVF) in the diffuse boundaries image, whereas it remains steady in SAM. A comprehensive analysis performed on the entire benchmark dataset is presented in Table 4, confirming these findings.
@@ 5.3 Floe count, size distribution, and sea ice concentration across the entire sample
Data extracted from all 20 images in the sample are compared with the benchmark manual segmentation, as illustrated in Fig. 9. The relative error (RE) in estimating floe numbers is presented in Fig. 9a. Among the methods evaluated, SAM-P performs the best, slightly underestimating the number of floes by approximately 2%, which highlights its capability to detect floes across images with varying clarity. In contrast, SAM-A and GVF are the least accurate, underestimating the manual count by approximately 28% and 26%, respectively. This reflects the difficulty in detecting floes in complex scenarios, particularly when dealing with unclear or irregular ice formations (cf. Figs. 7 and 8). For SAM-A, the automated initialization is a key limitation, as it prioritizes larger, more distinct floes while often overlooking smaller or poorly defined ones. The combined SAM-GVF approach mitigates some of the uncertainties in floe detection observed in both SAM-A and GVF, leading to a lower underestimation of the manual count (approximately 12%). This aligns with its performance on the distinct edges image, suggesting that, like SAM-P, it remains relatively consistent across varying image conditions.
Figure 9: Comparison of floe size statistics derived from segmentation methods GVF (green), SAM-A (red), SAM-P (pink), and SAM-GVF (purple) against the manual benchmark (orange). (a) Relative error (RE) in the number of detected floes compared to the manual benchmark (horizontal orange line). b) Violin plots of floe size distributions for each method, with the manual benchmark on the left. Squares inside violins represent medians; darker shaded areas indicate quartiles. (c,d,e,f) $Q$ plots comparing deciles (10% to 90%) of floe size distributions from segmentation methods against the manual benchmark. The solid line indicates perfect alignment. (g) Relative error (RE) in ice concentration compared to the manual benchmark (horizontal orange line).
@@ 6 Conclusion
The GVF method provided a reasonable floe size distribution due to its ability to capture floe shapes but was less effective in detecting all floes, impacting its accuracy in estimating ice concentration. SAM-A performed well on images with distinct floe edges, though its effectiveness decreased significantly in challenging environments, detecting fewer floes and providing the most divergent size distributions from the benchmark. SAM-A provided an accurate sea ice concentration estimate, albeit due to compensating factors: fewer, but larger detected floes. SAM-P was more effective at detecting floes, consistently producing accurate size distributions and robust sea ice concentration estimates across the entire sample. The SAM-GVF combination slightly improved upon SAM-P by reducing misinterpretations of complex floes, although overall performance remained similar, closely aligning with the benchmark in both floe size distribution and sea ice concentration.
SAM-P and SAM-GVF emerged as the most effective methods, demonstrating strong performance across multiple metrics, especially in accurately estimating sea ice concentration and improving floe detection. SAM-P’s computational efficiency makes it suitable for real-time applications requiring immediate results and for processing large datasets to derive statistical properties of sea ice. Conversely, SAM-GVF and GVF alone, despite their longer computational times, are better suited for detailed analyses where precise floe shape information is essential.
This study establishes robust methodological foundations for obtaining reliable quantitative data on Antarctic sea ice floe sizes in the MIZ from field-acquired imagery, addressing the scarcity and limitations inherent in direct field observations. Although SAM, as a generalist segmentation model, is not currently optimized for sea ice, this presents an opportunity for further training specifically for Antarctic sea ice conditions using our extensive dataset. Building on the demonstrated effectiveness of these segmentation methods, future research will extend the analysis to include the complete image collection from numerous Antarctic expeditions and various seasons, enabling comprehensive monitoring of the evolution of Antarctic sea ice.
@@
The 20 sample images used in this study, along with the image processing scripts, are available on GitHub (Passerotti, 2025).
@@ Acknowledgments
Appendix B Benchmark dataset photos and segmentation performance
@@
Figures 10 and 11 present the twenty photos comprising the benchmark dataset, illustrating a range of ice conditions and lighting environments. Approximately 30% of the dataset represents optimal conditions, featuring distinct ice floes with clearly defined edges, while the remaining 70% captures more complex scenarios characterized by diffuse sea ice boundaries.
The segmentation performance of the four methods applied to the entire benchmark dataset is summarized in Table 4, categorized by missing floes, misinterpreted floes, and hallucinations. Manual segmentation identifies 2,874 floes in total, with an average diameter of 3.3 m and an average ice concentration of 45%.
Benchmark dataset
Method Missing Misinterpreted Hallucinations
GVF 1030 2.09 m 5% 411 3.53 m 7% 212 1.96 m 1%
SAM-A 767 1.65 m 2% 235 5.36 m 8% 105 3.42 m 1%
SAM-P 578 1.98 m 2% 228 4.45 m 5% 658 1.32 m 1%
SAM-GVF 707 1.97 m 3% 166 3.98 m 3% 389 1.64 m 1%
The GVF method exhibits the highest number of missed floes (1,030), accounting for approximately 35% of the total. SAM-P, in contrast, misses the fewest floes ($\approx$20%), while SAM-A and SAM-GVF each miss slightly more than 25%. Across all methods, missed floes have an average diameter of around 2 m, covering between 2% and 5% of the benchmark image area. Misinterpretations range from 5% to 15% of the total floes, with SAM-GVF performing best (166 misinterpretations) and GVF performing worst (411 misinterpretations). These misinterpreted floes typically have larger average diameters than the missed floes, indicating these errors mostly result from merging adjacent floes into larger objects. Hallucinations vary significantly among the methods but generally affect only around 1% of the total image area, making their impact marginal in the overall assessment. The average diameter of hallucinated floes is typically below 2 m, except for SAM-A, which produces hallucinations averaging 3.4 m.
Overall, these results align closely with the findings from the two sample images presented in §5.2. The higher variability in hallucinations, particularly notable for SAM-P and SAM-GVF, primarily arises from image “22-07-24 02-52-21” (Fig. 11), where snow is misclassified as small floes, substantially increasing the hallucination count.

### v5

@@ 5.3.2 Diffuse sea ice boundaries
Of the various methods, SAM-P provides the best overall balance of accuracy and computational efficiency across both the distinct edges and diffuse boundaries images. SAM-P has the fewest missed floes and maintains stable, though slightly higher, rates of misinterpretations and hallucinations compared to SAM-GVF. While SAM-GVF performs marginally better in terms of misinterpretation and hallucination metrics, these improvements come at a substantial cost in computational time ($\approx$15 times slower per floe). SAM-A and GVF show significant deterioration, particularly with a substantial increase in missed floes. Additionally, computational time per floe increases for GVF (and consequently for SAM-GVF) in the diffuse boundaries image, whereas it remains steady in SAM.
5.3.3 Round circular ice floes
The manual segmentation identifies a total of 145 floes, averaging 3.31 m in diameter with an ice concentration of 31%. Using the GVF method, 24 floes are missed, four are misinterpreted, and four hallucinations are produced, collectively accounting for approximately 22% of floes not correctly identified (Fig. 9c). Missed floes have an average diameter of 1.5 m, misinterpretations average 3.5 m, and hallucinations average 0.8 m. Together, these errors account for less than 3% of the total surface area. Most misinterpretations arise from partial segmentation of larger floes, while hallucinations and misses involve smaller floes, indicating that GVF has difficulty detecting fine-scale features. Nevertheless, these errors minimally impact the final results, particularly when averaging floe diameters across the entire image.
Table 4: Segmentation results for the round circular floes (“22-07-24 11-17-13”) image using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Round circular floes image
GVF 24 1.51 m 1% 4 3.45 m 1% 4 0.80 m <1% 35 s
SAM-A 19 1.66 m 1% 8 5.31 m 4% 5 2.06 m <1% 0.8 s
SAM-P 9 2.44 m 1% 9 4.41 m 3% 25 0.69 m <1% 0.3 s
SAM-GVF 19 2.13 m 2% 3 3.05 m <1% 14 0.73 m <1% 6 s
Figure 9: Visual comparison between manual and automated segmentation methods for an image with round circular ice floes from the benchmark dataset (“22-07-24 11-17-13” in Fig. 13). The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
SAM-A demonstrates performance similar to GVF in terms of error quantity, missing 19 floes, misinterpreting eight, and generating five hallucinations (also totaling approximately 22%). The missing floes are slightly larger (1.6 m average diameter) and cover a similar area to GVF. Misinterpretations exhibit a larger average diameter (5.3 m), primarily due to the merging of adjacent floes, biasing the size distribution toward larger values. However, their limited number ensures a minimal impact on the overall results. Hallucinations average 2.1 m in diameter, arising mainly from darker oceanic features mistaken as floes (Fig. 9e), yet they affect less than 1% of the total area.
SAM-P misses the fewest floes (only nine), with nine misinterpretations and notably higher hallucinations (25), totaling approximately 43% of incorrectly identified floes. Misinterpretations have a slightly smaller average diameter (4.4 m) compared to SAM-A, as the model correctly separates some previously merged floes (Fig. 9g). Missing floes tend to be larger (2.4 m), frequently occurring when one floe in a merged pair is correctly identified while the other is missed entirely. Hallucinations, though numerous, have a mean diameter of 0.7 m, primarily arising from image noise near larger floes, and cover less than 1% of the area. In total, all errors combined affect less than 4% of the image area, confirming SAM-P’s strong ability to accurately estimate floe characteristics and ice concentration.
SAM-GVF misses more floes than SAM-P (19 floes, similar to SAM-A and GVF), but has fewer misinterpretations (only three) and a small average misinterpretation diameter (3.05 m). Hallucinations are similar to SAM-P, averaging 0.7 m and covering less than 1% of the area. Combined, SAM-GVF’s errors represent 25% of the floes, consistent with SAM-A and GVF, but with minimal impact on ice concentration estimates.
Overall, the round circular floes image represents optimal conditions, featuring relatively few, well-spaced floes with uniform shape and size. Consequently, all methods perform well, particularly in capturing average floe diameters. Under such optimal conditions, computational efficiency becomes especially important: SAM-P is notably the fastest method (0.3 s per floe), SAM-A takes slightly more than twice as long, SAM-GVF is roughly twenty times slower, and GVF is the slowest, requiring approximately 35 s per floe.
5.3.4 Angular square ice floes
Segmentation performance slightly degrades in the angular square floes scenario due to the less regular and more heterogeneous floe shapes and sizes. For this sample image, manual segmentation identifies 127 floes with an average diameter of 3.5 m and an ice concentration of 49%. The GVF method misses 59 floes, more than double the number compared to the round circular floes case. The missed floes have an average diameter of 1.2 m, affecting approximately 2% of the total image area, confirming their limited impact on ice concentration but significant influence on the smaller size tail of the floe distribution. A substantial difference emerges in misinterpretations, totaling 38 (nearly tenfold higher), with an average diameter of 4 m, impacting 12% of the surface area. These errors predominantly result from partial segmentations of larger, irregularly shaped floes, indicating GVF’s difficulty in handling large floes formed by merged pancakes (Fig. 10c). Hallucinations remain minimal, consistent with the previous observations. Overall, nearly 80% of the floes are incorrectly identified.
Table 5: Segmentation results for the angular square floes image (“22-07-21 12-48-26”) using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Angular square floes image
GVF 59 1.20 m 2% 38 3.94 m 12% 2 0.97 m <1% 70 s
SAM-A 42 1.18 m 2% 17 4.59 m 6% 1 5.07 m <1% 0.8 s
SAM-P 52 1.08 m 1% 7 6.18 m 5% 2 0.50 m <1% 0.3 s
SAM-GVF 55 1.09 m 1% 7 4.53 m 3% 1 0.30 m <1% 6 s
Figure 10: Visual comparison between manual and automated segmentation methods for an image with angular square ice floes from the benchmark dataset (“22-07-21 12-48-26” in Fig. 12). The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
SAM-A exhibits comparable missed floes metrics to GVF, with 42 misses, double the round circular floes scenario. The missed floes’ average diameter is slightly smaller (1.2 m), impacting about 2% of the surface area. Misinterpretations also double to 17, averaging 4.6 m in diameter and covering 6% of the area. While SAM-A effectively identifies larger floes despite irregular shapes, segmentation remains imperfect due to pronounced textures causing partial segmentation and merged adjacent floes (Fig. 10e). Only one significant hallucination occurs, clearly identifiable as an erroneous large water gap classified as a floe.
SAM-P misses 52 floes, approximately five times more than in the round circular scenario, averaging 1.1 m and covering 1% of the image area. However, misinterpretations are fewer (only 7), with a larger average diameter (6.2 m), covering 5% of the area, demonstrating consistent segmentation capabilities regardless of floe shapes. Hallucinations are negligible in both number and impact.
SAM-GVF closely matches SAM-P performance, missing 55 floes (average diameter 1.1 m) and producing only one small hallucination (diameter 0.30 m) with no impact on overall sea ice characteristics. Misinterpretations remain at seven, but have smaller diameters averaging less than 4.5 m and cover only 3% of the image area, indicating improved accuracy in separating merged floes compared to SAM-P.
Overall, GVF shows notably higher errors compared to the SAM-based methods. In terms of surface area affected by errors, SAM-GVF performs best at 4%, followed by SAM-P (6%), SAM-A (8%), and GVF (14%). Computationally, SAM-P remains highly efficient at 0.3 s per floe, unaffected by floe complexity, whereas GVF’s computational time doubles to 70 s per floe. SAM-GVF maintains consistent computational efficiency at approximately 6 s per floe, benefiting from SAM’s speed.
@@ 5.4 Floe count, size distribution, and sea ice concentration across the entire sample
Figure 11: Comparison of floe size statistics derived from segmentation methods GVF (green), SAM-A (red), SAM-P (pink), and SAM-GVF (purple) against the manual benchmark (orange). (a) Relative error (RE) in the number of detected floes compared to the manual benchmark (horizontal orange line). b) Violin plots of floe size distributions for each method, with the manual benchmark on the left. Squares inside violins represent medians; darker shaded areas indicate quartiles. (c,d,e,f) $Q$ plots comparing deciles (10% to 90%) of floe size distributions from segmentation methods (D Mod.) against the manual benchmark (D Bmk.). The solid line indicates perfect alignment. (g) Relative error (RE) in ice concentration compared to the manual benchmark (horizontal orange line).
Data extracted from all 20 images in the sample are compared with the benchmark manual segmentation, as illustrated in Fig. 11. The relative error (RE) in estimating floe numbers is presented in Fig. 11a. Among the methods evaluated, SAM-P performs the best, slightly underestimating the number of floes by approximately 2%, which highlights its capability to detect floes across images with varying clarity. In contrast, SAM-A and GVF are the least accurate, underestimating the manual count by approximately 28% and 26%, respectively. This reflects the difficulty in detecting floes in complex scenarios, particularly when dealing with unclear or irregular ice formations (cf. Figs. 8 and 10). For SAM-A, the automated initialization is a key limitation, as it prioritizes larger, more distinct floes while often overlooking smaller or poorly defined ones. The combined SAM-GVF approach mitigates some of the uncertainties in floe detection observed in both SAM-A and GVF, leading to a lower underestimation of the manual count (approximately 12%). This aligns with its performance on the distinct edges image, suggesting that, like SAM-P, it remains relatively consistent across varying image conditions.
@@ 6 Conclusion
The GVF method provided a reasonable floe size distribution due to its ability to capture floe shapes, particularly when floes were regular, circular, and medium sized, but it was less effective in detecting all floes, impacting its accuracy in estimating ice concentration. SAM-A consistently performed well on images with distinct floe edges regardless of floe shape, although its effectiveness decreased significantly in challenging environments, detecting fewer floes and providing the most divergent size distributions from the benchmark. SAM-A provided an accurate sea ice concentration estimate, albeit due to compensating factors: fewer, but larger detected floes. SAM-P was more effective at detecting floes, consistently producing accurate size distributions and robust sea ice concentration estimates across the entire sample. The SAM-GVF combination slightly improved upon SAM-P by reducing misinterpretations of complex floes, although overall performance remained similar, closely aligning with the benchmark in both floe size distribution and sea ice concentration.
SAM-P and SAM-GVF emerged as the most effective methods, demonstrating strong performance across multiple metrics, especially in accurately estimating sea ice concentration and improving floe detection. SAM-P’s computational efficiency makes it suitable for real-time applications requiring immediate results and for processing large datasets to derive statistical properties of sea ice. Conversely, SAM-GVF, despite its longer computational times, is better suited for detailed analyses where precise floe shape information is essential.
This study establishes robust methodological foundations for obtaining reliable quantitative data on sea ice floes in the MIZ from field-acquired imagery. Due to logistical and observational constraints, Antarctic sea ice datasets, especially floe size data, remain limited. By developing an effective segmentation methodology, we facilitate improved utilization of field imagery, increasing the quantity and quality of data obtained from these observational campaigns. Although SAM, as a generalist model, is not currently optimized for sea ice, our extensive dataset provides an opportunity for further training specifically for Antarctic conditions. Future research will extend this analysis to the complete image collection from multiple Antarctic expeditions and seasons, enabling comprehensive monitoring of the evolution of Antarctic sea ice.
@@
The 20 sample images used in this study, along with the image processing scripts, are available on GitHub Passerotti (2025).
@@ Acknowledgments
Appendix B Benchmark dataset photos
@@
Figures 12 and 13 present the twenty photos comprising the benchmark dataset, illustrating a range of ice conditions and lighting environments. Approximately 30% of the dataset represents optimal conditions, featuring distinct ice floes with clearly defined edges, while the remaining 70% captures more complex scenarios characterized by diffuse sea ice boundaries. Additionally, roughly half of the benchmark dataset consists of floes with round or circular shapes, while the other half features floes that are more angular, squared, and less regular in shape.